# Disorder-induced strong-field strong-localization in 2D systems

### v1

@@ III Impurity density Consideration
Our SCBA treatment of the disorder broadening leading to Eqs. (11) and (12) for the critical filling for the localization transition explicitly assumes short-range disorder, and the zero-field disorder is proportional to the impurity density $n_{i}$ in the leading order theory:
@@
where $a_{B}$ is the effective Bohr radius and $\kappa$ is the lattice dielectric constant for the semiconductor. The screened Coulomb disorder potential is then given by:
@@
which is independent of the carrier density and $q$, reflecting the energy independence of the 2D density of states. Thus, the 2D screened Coulomb disorder is effectively short-ranged. Using the Born approximation to calculate the impurity scattering induced zero-field broadening then gives:
@@
This scaling is linear instead of being a square root in $\Gamma_{0}/\hbar\omega_{c}$ as in Eq. (7), but they have the same qualitative dependence: both predict a critical filling increasing with increasing (decreasing) with zero-field broadening (cyclotron frequency). This is indeed the qualitative behavior of the observed $\nu_{c}$ in the transition to the strongly insulating state at low LLL filling, which is accepted uncritically as the signature for a transition to a strong-field WC. In our non-interacting theory, however, the transition is to a strongly localized Anderson insulator.
@@ IV Phase Diagram
We next discuss the qualitative phase diagram implied by the above considerations, now putting in empirically and phenomenologically the known 2D strong-field interaction effects. Much work has been done on the possible phase diagram in the LLL in the absence of disorder, and there are experimental claims of quantitative agreements with such theoretical phase diagrams focusing on the clean system (“no disorder”) where the competing phases are WC and FQHL [39, 40, 41]. The problem is that the theoretical calculations uncritically ignore disorder and compare ground state energies using model wave functions for WC and FQHL. On the other hand, the experimental phase diagram is crucially dependent on disorder since the key experiment is the observation of a strongly resistive insulating state at low LLL filling which is automatically assumed to be the onset of the WC phase overcoming the FQHL phase. Ref. 1 casts doubt on this interpretation since the WC seems to be generically present everywhere, becoming an amorphous solid at low LLL filling, where presumably the corresponding transport manifests the strongly insulating state.
@@
Combining the existing theoretical comparisons between WC and FQHL energetics with our disorder results, we propose a schematic phase diagram shown in Figs. 3 and 4, each highlighting schematically and qualitatively different aspects of the competing phases arising from the existence of disorder as well as WC, IQHE, FQHE, which are all distinct phases arising from different physics. The $\Gamma=E_{F}$ line is the key disorder contribution emphasizing the possible disorder-induced Anderson insulator phase.
@@
The broad phase diagram (in color) shown in Fig. 3 is a high-level construct combining our earlier work on IQHE [31] with the current work. It shows the broad features arising from IQHE and FQHE coupled with disorder, with the FQHE part being associated with interactions. For weak disorder ($\Gamma<T$), thermal activation delocalizes all localized states (shaded red in Fig. 3) and completely suppresses the quantum Hall effect, so there is no plateau in $\sigma_{xy}=\nu e^{2}/h$ and $\sigma_{xx}$ becomes a constant. In the intermediate disorder regime ($T<\Gamma<\hbar\omega_{c}$), the broadening of the extended states in the filling factor follows $T/\Gamma$. Because Fig. 3 employs the inverse filling factor as the horizontal axis, where the interval between consecutive integer fillings $\nu$ and $\nu+1$ scales as $\sim 1/\nu$, the corresponding broadening in this inverse filling factor axis follows $T/(\Gamma\nu)$. When the disorder strength increases to $\hbar\omega_{c}<\Gamma<N\hbar\omega_{c}\sim E_{F}$, this broadening in filling factor saturates at $T/\hbar\omega_{c}$ [or $T/(\hbar\omega_{c}\nu)$ in the inverse filling factor as shown in Fig. 3]. For strong disorder ($\Gamma>N\hbar\omega_{c}$), the extended states shift to higher filling factors. At sufficiently large filling factors, where the cyclotron radius $R_{c}$ is larger than the disorder potential correlation length $d$ so that $R_{c}/d>\max(1,\hbar\omega_{c}/\Gamma)$, the diffusion of electron guiding centers becomes efficient, disrupting semiclassical localization in the quantum Hall regime [31]. As a result, the system undergoes a crossover from IQHE phase to the no-QHE metallic phase where $\sigma_{xy}\propto 1/B$ and $\sigma_{xx}$ constant in $B$ follows the conventional Drude formula. If $T>\hbar\omega_{c}$, all QHE phases are destroyed and leading to the semiclassical metallic phase without QHE plateau in $\sigma_{xy}$.
@@
The most relevant phase diagram for the current work is in Fig. 4, which restricts to the LLL, and shows qualitatively all the relevant phases: FQHL, WC, AS. Below the $\Gamma=E_{F}$ line away from $\nu=1/2$, regime in between two FQH states is generically occupied by WC (or a similar interacting phase, e.g., stripes, with no topological order and no FQHE), as shown in Figs. 2 and 3 of Ref. 1. Figure 4 conjectures that in the limit of zero disorder, the FQHE is restricted to a set of measure zero occurring precisely at odd-denominator filling factors. Both WC and FQHL are truncated by the disorder line, and if the disorder is strong, then neither would manifest itself (which is the situation in all samples prior to the 1982 discovery of FQHE). If the disorder is weak, however, many FQHE at many odd denominators would appear, which is the current situation in the very best samples (both GaAs and graphene) where FQHL exists experimentally to denominators as large as 13 [44, 16]. But eventually, the disorder line comes down at low filling, and may touch the WC/FQHL sequence, terminating their existence with a crossover to the AS phase. Details would depend not only on the disorder strength, but also on the temperature. Other than a general statement that in high (low) –quality samples many (few or none) FQHL phases would be observed (Ref. 1 manifests only one at 1/3), very little more can be stated generically. Note that the essence of this phase diagram is that the WC in the presence of disorder goes continuously from a pinned WC to a completely random and amorphous AS with little spatial order depending on both the filling factor and the disorder strength.
@@ V Conclusion
Our prediction of $\nu_{c}\sim(\Gamma_{0}/\hbar\omega_{c})^{1/2}$ or $(\Gamma_{0}m/n)$—the two formulae are equivalent since $\nu=2\pi l_{B}^{2}n$—is in principle verifiable, but in general the disorder-broadening is unknown. Also, our theory includes only the physics of Anderson localization, which would surely be modified by the presence of interaction in the problem. Predicting $\nu_{c}$ accurately in realistic samples including disorder and interaction microscopically (and allowing the possibility for all possible quantum phases—FQHL, WC, metal, AS, …) is completely out of scope for any current theoretical tools. We believe that the theories based just on energetics without including disorder miss an essential ingredient of the experiments in real samples. We do emphasize that our purely disorder-based simplified theory predicts the correct experimental trends in $\nu_{c}$, namely, the decrease in $\nu_{c}$ with increasing mobility and decreasing magnetic field and density.

### v2

@@ III Impurity density Consideration
Our SCBA treatment of disorder broadening, which yields the critical filling factors in Eqs. (11) and (12), explicitly assumes short-range disorder. In this leading-order theory, the zero-field broadening is directly proportional to the impurity density $n_{i}$:
@@
where $a_{B}$ is the effective Bohr radius and $\kappa$ is the lattice dielectric constant for the semiconductor. The Fourier transformation of the screened Coulomb disorder potential is then given by:
@@
which is independent of the carrier density and momentum transfer $q$, reflecting the energy independence of the 2D density of states. Thus, the 2D screened Coulomb disorder is effectively short-ranged. Using the Born approximation to calculate the impurity scattering induced zero-field broadening then gives:
@@
This scaling is linear instead of being a square root in $\Gamma_{0}/\hbar\omega_{c}$ as in Eq. (7), but they have the same qualitative dependence: both predict a critical filling increasing with increasing (decreasing) zero-field broadening (cyclotron frequency). This is indeed the qualitative behavior of the observed $\nu_{c}$ in the transition to the strongly insulating state at low LLL filling, which is accepted uncritically as the signature for a transition to a strong-field WC. In our non-interacting theory, however, the transition is to a strongly localized Anderson insulator.
@@ IV Phase Diagram
We next discuss the qualitative phase diagram implied by the above considerations, now putting in empirically and phenomenologically the known 2D strong-field interaction effects. Extensive research has explored the LLL phase diagram in the limit of zero disorder, with several experimental studies claiming quantitative agreement with these ”clean” theoretical models [42, 43, 44]. However, these theoretical treatments often neglect disorder entirely, comparing ground-state energies only for WC and FQHL wavefunctions. In contrast, experimental phase diagrams are inherently sensitive to disorder. The standard interpretation identifies the onset of a strongly resistive state at low filling as the WC phase overcoming the FQH phase. However, this interpretation often overlooks the possibility of a disorder-induced AS. Crucially, both a pinned WC and an AS can exhibit strongly insulating behavior, making transport data alone insufficient to distinguish between them.
@@
Combining the existing theoretical comparisons between WC and FQHL energetics with our disorder results, we propose a schematic phase diagram shown in Figs. 3 and 4, each highlighting schematically and qualitatively different aspects of the competing phases arising from the existence of disorder as well as WC, IQHE, FQHE, which are all distinct phases arising from different physics. The $\Gamma=E_{F}$ line is the key disorder contribution emphasizing the possible disorder-induced insulating AS phase.
@@
The broad phase diagram (in color) shown in Fig. 3 is a high-level construct combining our earlier work on IQHE [34] with the current work. It shows the broad features arising from IQHE and FQHE coupled with disorder, with the FQHE part being associated with interactions. For weak disorder ($\Gamma<T$), thermal activation delocalizes all states originally localized at $T=0$ (shaded red in Fig. 3) and completely suppresses the QHE, so there is no plateau in $\sigma_{xy}=\nu e^{2}/h$. In the intermediate disorder regime ($T<\Gamma<\hbar\omega_{c}$), the broadening of the extended states in the filling factor follows $T/\Gamma$. Because Fig. 3 employs the inverse filling factor as the horizontal axis, where the interval between consecutive integer fillings $\nu$ and $\nu+1$ scales as $\sim 1/\nu$, the corresponding broadening in this inverse filling factor axis follows $T/(\Gamma\nu)$. When the disorder strength increases to $\hbar\omega_{c}<\Gamma<N\hbar\omega_{c}\sim E_{F}$, this broadening in filling factor saturates at $T/\hbar\omega_{c}$ [or $T/(\hbar\omega_{c}\nu)$ in the inverse filling factor as shown in Fig. 3]. For strong disorder ($\Gamma>N\hbar\omega_{c}$), the extended states shift to higher filling factors, which is known as the floating physics. At sufficiently large filling factors, where the cyclotron radius $R_{c}$ is larger than the disorder potential correlation length $d$ so that $R_{c}/d>\max(1,\hbar\omega_{c}/\Gamma)$, the diffusion of electron guiding centers becomes efficient, disrupting semiclassical localization in the quantum Hall regime [34]. As a result, the system undergoes a crossover from IQHE phase to the no-QHE metallic phase where $\sigma_{xy}\propto 1/B$ and $\sigma_{xx}$ constant in $B$ follows the conventional Drude formula. If $T>\hbar\omega_{c}$, all QHE phases are destroyed and leading to the semiclassical metallic phase without QHE plateau in $\sigma_{xy}$.
@@
The most relevant phase diagram for the current work is in Fig. 4, which restricts to the LLL, and shows qualitatively all the relevant phases: FQHL, WC, AS. Below the $\Gamma=E_{F}$ line away from $\nu=1/2$, regime in between two FQH states is generically occupied by WC (or a similar interacting phase, e.g., stripes, with no topological order and no FQHE), as shown in Figs. 2 and 3 of Ref. 1. Figure 4 conjectures that in the limit of zero disorder, the FQHE is restricted to a set of measure zero occurring precisely at odd-denominator filling factors. Both WC and FQHL are truncated by the disorder line, and if the disorder is strong, then neither would manifest itself (which is the situation in all samples prior to the 1982 discovery of FQHE). If the disorder is weak, however, many FQHE at many odd denominators would appear, which is the current situation in the very best samples (both GaAs and graphene) where FQHL exists experimentally to denominators as large as 13 [47, 16]. But eventually, the disorder line comes down at low filling, and may touch the WC/FQHL sequence, terminating their existence with a crossover to the AS phase. Details would depend not only on the disorder strength, but also on the temperature. Beyond the general observation that high-quality samples exhibit numerous FQH states while lower-quality ones show few or none (with Ref. 1 manifesting only the $1/3$ and $2/5$ states) little else can be stated generically. Note that the essence of this phase diagram is that the WC in the presence of disorder goes continuously from a pinned WC to a completely random and amorphous AS with little spatial order depending on both the filling factor and the disorder strength.
@@ V Conclusion
Our prediction of $\nu_{c}\sim(\Gamma_{0}/\hbar\omega_{c})^{1/2}$ or $(\Gamma_{0}m/\hbar^{2}n)$–the two formulae are equivalent since $\nu=2\pi l_{B}^{2}n$–is in principle verifiable, but in general the disorder-broadening is unknown. Also, our theory includes only the physics of Anderson localization, which would surely be modified by the presence of interaction in the problem. Predicting $\nu_{c}$ accurately in realistic samples including disorder and interaction microscopically (and allowing the possibility for all possible quantum phases—FQHL, WC, AS, …) is completely out of scope for any current theoretical tools. We believe that the theories based just on energetics without including disorder miss an essential ingredient of the experiments in real samples. We do emphasize that our purely disorder-based simplified theory predicts the correct experimental trends in $\nu_{c}$, namely, the decrease in $\nu_{c}$ with increasing mobility.
We emphasize that both interaction and disorder are important in the LLL physics, as is well-known and also is obvious from the appearance of both FQHE and WC in the results of Ref. [1]. We are by no means saying that interaction can be ignored. We are, however, claiming that the uncritical assertion of associating the strongly insulating state at low LLL filling, often observed in transport measurements, automatically with a WC formation may have to be re-evaluated in view of the results of Ref. [1]. We believe that the formation of the strongly insulating state has a lot to do with the amount of disorder in the system associated with the inevitable existence of random charged impurities even in the best 2D samples. In particular, Imry-Ma arguments [56] rule out the existence of long range order in the presence of quenched random impurities which act as a random field. Therefore, the WC formation in the strict sense is not allowed. In fact, there is no spontaneous symmetry breaking here at all since the starting Hamiltonian with random charged impurities does not have any translational order already– translational symmetry is broken explicitly by the random field associated with disorder and not spontaneously by interaction. The only question is the extent of the short range order associated with the local (not global) breaking of the rotational symmetry apparent in Fig. 2 of Ref. 1 for $\nu>0.093$. Since eventually the local hexagonal pattern disappears with the structure factor manifesting only a ring for $\nu=\nu_{c}\sim 0.093$ in Ref. 1, it is reasonable to assume that the system is an amorphous AS for $\nu<\nu_{c}$ instead of being any type of Wigner solid. Of course, a highly fragmented (through strong pinning by many random impurities) strongly pinned Wigner crystal with no spatial ordering is the same as the amorphous AS, and it may then simply be semantics to call it AS or strongly pinned WS. All we are saying is that it may be misleading to emphasize only the WC aspect of the insulating phase over the AS aspect, particularly since the best theoretical estimates seem to indicate that in a pristine system, the FQHL may remain the ground state for arbitrarily low LLL filling in the absence of any disorder [30, 53].